If f(x)= x^2 -1 and g (x) =2x +1, show that f[g(x)]=4x(x+1).

1
Expert's answer
2022-06-07T13:09:59-0400

Given:

"f(x)=x^2-1,\quad g(x)=2x+1"



"f(g(x))=(2x+1)^2-1=4x^2+4x+1-1\\
=4x^2+4x=4x(x+1)"